OpenStudy (anonymous):

\[2 \int \int y(t)dtdt+\int y(t) \, dt=\sin (x)\]

OpenStudy (dumbcow):

where does the x come from? should it be sin(t) are you supposed to find y(t)

OpenStudy (anonymous):

sorry, \[2 \int \int y(t)dtdt+\int y(t) \, dt=\sin (t)\]

OpenStudy (anonymous):

Where are there two dts?

OpenStudy (anonymous):

Err why rather.

OpenStudy (anonymous):

I mean I suppose it's fine to double integrate, but I've never seen that before..

OpenStudy (dumbcow):

im guessing its like the inverse of the 2nd derivative let u(t) by the result, then u'' = y(t)

OpenStudy (anonymous):

double integral with same variable , I believe

OpenStudy (anonymous):

Or sorry rather you're saying that u'(t) + 2u(t) = sin(t) where u''(t) = y(t)

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

So solve the differential for u(t), then differentiate it twice to find y?

OpenStudy (anonymous):

Ok, so it's just linear differential from here

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thank you much to both of you